? acidic V basic ? amphoteric So3: acidic basic amphoteric Write a balanced molecular equation, including physical states, for the reaction of each oxide with excess water Bao(s ) + H 2 0(1) ?? ? SO 3 (g)+ H 200)-
Answer #1

BaO(s) + H2O(l) = Ba(OH)2(aq)

SO3(g) + H2O(l) = H2SO4(aq)

If you satisfied with the solution please like it thankxx

If you have any query please comment